The line above goes through points (-5,5) and (0,1). Which equation matches the line plot above?
Answer by stanbon(75887)

The line above goes through points (-5,5) and (0,1).
Which equation matches the line plot above?
--------
slope = (1-5)/(0--5) = -4/5
--------
y-intercept:
f(0) = 1
================
y = (-4/5)x
================
Cheers,
Stan H.
================